OpenStudy (lemanual):

Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. : x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive? Do not enter comma for placeholder values.

OpenStudy (anonymous):

x = number of votes Maya received: this is what you want to solve for y = total number of votes cast y is given to you = 125,000 Solve for x by plugging in y: x - 62,500 = 0.25(125,000) x - 62,500 = 31,250 x = 93,750 = Number of votes Maya received

OpenStudy (anonymous):

Nice salute there son ;) I take it you want to join the ranks in the future! Awesome. This is really easy, you just need to plug and chug the numbers. x = number of votes Maya received: this is what you want to solve for y = total number of votes cast 125,000=y so: x-62500= .25*125,000 x-62,500=31,250 x=31,250+62,500 x=93, 750 You can also double check the math and see .75 (Maya's portion of the votes)*125,000 (total number of votes)=93,750.

OpenStudy (lemanual):

thnx both if you